OpenStudy (anonymous):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = ?, b = 11, c = 17 A. 12.75 B. 13.20 C. 13.12 D. 12.96 @REDNECK4302

OpenStudy (dmb1215):

do you know the Pythagorean theorem

OpenStudy (anonymous):

Not really, I don't know anything about it really, it confuses me. I'm home schooled so I've been just skipping it until I HAVE to do it. Which was a REALLY stupid idea.

OpenStudy (dmb1215):

d

OpenStudy (anonymous):

A SQUARED PLUS B SQUARED = C SQUARED

OpenStudy (anonymous):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = 16, b = 28, c = ? A. 31.46 B. 31.94 C. 32.09 D. 32.25

jimthompson5910 (jim_thompson5910):

Use the Pythagorean Theorem. Plug in b = 11, c = 17 and solve for 'a' to get \[\Large a^2 + b^2 = c^2\] \[\Large a^2 + 11^2 = 17^2\] \[\Large a^2 + 121 = 289\] \[\Large a^2 = 289-121\] \[\Large a^2 = 168\] \[\Large a = \sqrt{168}\] \[\Large a \approx 12.96148 \ ... \ \text{ Use a calculator here}\] \[\Large a \approx 12.96 \ ... \ \text{ Round to 2 decimal places}\]

jimthompson5910 (jim_thompson5910):

For the next one, you'll have... \[\Large a^2 + b^2 = c^2\] \[\Large 16^2 + 28^2 = c^2\] and I'll let you finish up